# Promotion of Scientific Publications on ArXiv and X Is on the Rise and Impacts Citations

### v1

@@ A. Distribution of all citations for each treatment
Figure 8: Box plots for the number of citations for each treatment. The red line in the middle of the box represents the median and the orange triangle represents the mean for (a) Number of citations for all papers, (b) ArXiv-first, (c) Revision, (d) Tweeted, and (e) Tweeted by author.
B. Missing top conferences from WikiCFP
@@ D. Name matching algorithm to find tweets from authors
Treatment Year TMLE Standardization
ATE p-value ATE p-value
First $\mathbf{0.5\pm 0.5}$ $\mathbf{0.039}$ $\mathbf{0.3\pm 0.1}$ $\mathbf{<10^{-6}}$
Arxiv-first Second $\mathbf{2.7\pm 1.7}$ $\mathbf{0.001}$ $\mathbf{1.3\pm 0.4}$ $\mathbf{<10^{-6}}$
effect Third $\mathbf{4.7\pm 3.6}$ $\mathbf{0.010}$ $\mathbf{2.2\pm 0.8}$ $\mathbf{<10^{-6}}$
Fourth $6\pm 6.1$ $0.051$ $\mathbf{2.7\pm 1.4}$ $\mathbf{5.1\times 10^{-5}}$
Fifth $\mathbf{7.2\pm 6}$ $\mathbf{0.017}$ $\mathbf{3\pm 1.4}$ $\mathbf{2.7\times 10^{-5}}$
First $\mathbf{0.5\pm 0.4}$ $\bf{0.040}$ $\mathbf{0.3\pm 0.1}$ $\mathbf{<10^{-6}}$
Revision Second $\mathbf{2.5\pm 1.4}$ $\mathbf{0.008}$ $\mathbf{1.5\pm 0.4}$ $\mathbf{<10^{-6}}$
effect Third 4.5 $\pm$ 3.3 $\mathbf{0.008}$ $\mathbf{2.9\pm 0.8}$ $\mathbf{<10^{-6}}$
Fourth $5.6\pm 5.7$ $0.056$ $\mathbf{3.9\pm 1}$ $\mathbf{<10^{-6}}$
Fifth $4.9\pm 7.4$ $0.194$ $\mathbf{3.8\pm 1.3}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.5\pm 0.2}$ $\mathbf{<10^{-6}}$ $\mathbf{0.6\pm 0.1}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{4.9\pm 0.7}$ $\mathbf{<10^{-6}}$ $\mathbf{1.6\pm 0.3}$ $\mathbf{<10^{-6}}$
tweeting Third $\mathbf{9.9\pm 1.6}$ $\mathbf{<10^{-6}}$ $\mathbf{2.9\pm 0.4}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{13.9\pm 2.5}$ $\mathbf{<10^{-6}}$ $\mathbf{3.5\pm 0.6}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{14.0\pm 3.2}$ $\mathbf{<10^{-6}}$ $\mathbf{3.0\pm 0.7}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.1\pm 0.3}$ $\mathbf{<10^{-6}}$ $\mathbf{0.6\pm 0.2}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{1.1\pm 1}$ $\mathbf{0.037}$ $\mathbf{1.7\pm 0.6}$ $\mathbf{<10^{-6}}$
author(s) Third $1.9\pm 2.03$ $0.072$ $\mathbf{3.1\pm 1.8}$ $\mathbf{5\times 10^{-4}}$
tweeting Fourth $\mathbf{4.3\pm 3.2}$ $\mathbf{0.008}$ $\mathbf{4.1\pm 3.1}$ $\mathbf{0.009}$
Fifth $\mathbf{13.5\pm 5.5}$ $\mathbf{<10^{-6}}$ $\mathbf{5.3\pm 4.33}$ $\mathbf{0.015}$
@@
E. Average treatment effects for the first through fifth year of publication
@@
Treatment Year TMLE Standardization
ATE p-value ATE p-value
First $0.09\pm 0.10$ $0.108$ $\mathbf{0.05\pm 0.03}$ $\mathbf{7.5\times 10^{-5}}$
Arxiv-first Second $\mathbf{0.49\pm 0.36}$ $\mathbf{0.008}$ $\mathbf{0.21\pm 0.08}$ $\mathbf{<10^{-6}}$
effect Third $0.59\pm 0.90$ $0.196$ $\mathbf{0.33\pm 0.21}$ $\mathbf{0.002}$
Fourth $\mathbf{0.98\pm 0.96}$ $\mathbf{0.046}$ $\mathbf{0.38\pm 0.22}$ $\mathbf{0.001}$
Fifth $\mathbf{1.10\pm 0.88}$ $\mathbf{0.015}$ $\mathbf{0.38\pm 0.20}$ $\mathbf{2\times 10^{-4}}$
First $0.04\pm 0.10$ $0.426$ $\mathbf{0.05\pm 0.02}$ $\mathbf{4\times 10^{-6}}$
Revision Second $\mathbf{0.46\pm 0.33}$ $\mathbf{0.005}$ $\mathbf{0.28\pm 0.08}$ $\mathbf{<10^{-6}}$
effect Third 0.73 $\pm$ 78 $0.063$ $\mathbf{0.48\pm 0.15}$ $\mathbf{<10^{-6}}$
Fourth $0.76\pm 1.28$ $0.244$ $\mathbf{0.58\pm 0.20}$ $\mathbf{<10^{-6}}$
Fifth $0.55\pm 1.60$ $0.501$ $\mathbf{0.54\pm 0.27}$ $\mathbf{7.1\times 10^{-5}}$
First $\mathbf{0.32\pm 0.05}$ $\mathbf{<10^{-6}}$ $\mathbf{0.11\pm 0.02}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{1.01\pm 0.16}$ $\mathbf{<10^{-6}}$ $\mathbf{0.30\pm 0.05}$ $\mathbf{<10^{-6}}$
tweeting Third $\mathbf{1.79\pm 0.33}$ $\mathbf{<10^{-6}}$ $\mathbf{0.49\pm 0.09}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{2.20\pm 0.50}$ $\mathbf{<10^{-6}}$ $\mathbf{0.55\pm 0.12}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{2.11\pm 0.61}$ $\mathbf{<10^{-6}}$ $\mathbf{0.47\pm 0.11}$ $\mathbf{<10^{-6}}$
First $\mathbf{0.18\pm 0.05}$ $\mathbf{3.2\times 10^{-6}}$ $\mathbf{0.13\pm 0.06}$ $\mathbf{2\times<10^{-5}}$
Effect of Second $\mathbf{0.63\pm 0.27}$ $\mathbf{6.8\times 10^{-6}}$ $\mathbf{0.32\pm 0.18}$ $\mathbf{3\times 10^{-4}}$
author(s) Third $0.31\pm 0.47$ $0.187$ $\mathbf{0.61,\pm 0.44}$ $\mathbf{0.007}$
tweeting Fourth $\mathbf{2.43\pm 1.15}$ $\mathbf{3.2\times 10^{-5}}$ ${0.79\pm 0.85}$ 0.0710.071{0.071}0.071
Fifth $\mathbf{1.74\pm 0.80}$ $\mathbf{2\times 10^{-5}}$ ${0.94\pm 0.98}$ 0.0590.059{0.059}0.059

### v2

@@ A. Distribution of all citations for each treatment
Figure 8: Box plots for the number of citations for each treatment for computer science. The red line in the middle of the box represents the median and the orange triangle represents the mean for (a) Number of citations for all papers, (b) ArXiv-first, (c) Revision, (d) Tweeted, and (e) Tweeted by author.
Figure 9: Box plots for the number of citations for each treatment for physics. The red line in the middle of the box represents the median and the orange triangle represents the mean for (a) Number of citations for all papers, (b) ArXiv-first, (c) Revision, (d) Tweeted, and (e) Tweeted by author.
B. Missing top conferences and dates from WikiCFP
We gathered all conferences from CSRankings.org and found that 8 of the top conferences were missing from WikiCFP which we added to our dataset(listed below). In addition to these 8 conferences, we identified around 400 conference editions with missing or incomplete WikiCFP information. We annotated their dates manually, using information publicly available on the internet, to include as many important computer science venues as possible and to increase our dataset size.
@@ D. Name matching algorithm to find tweets from authors
Outcome Treatment   Computer Science Physics
Standardization Relative p-value Train Cross-val Standardization Relative p-value Train Cross-val
ATE ATE $R^{2}$ $R^{2}$ ATE ATE $R^{2}$ $R^{2}$
All citations ArXiv-first effect $9.0\pm 3.7$ 0.13 $<10^{-6}$ 0.49 0.19 $0.9\pm 0.3$ 0.05 $<10^{-6}$ 0.51 0.17
Revision Effect $12.0\pm 3.6$ 0.17 $<10^{-6}$ 0.48 0.14 $3.0\pm 0.4$ 0.16 $<10^{-6}$ 0.52 0.17
Effect of Tweeting $11.9\pm 1.9$ 0.16 $<10^{-6}$ 0.62 0.26 $2.0\pm 0.2$ 0.08 $<10^{-6}$ 0.58 0.24
Effect of Author(s) $14.8\pm 8.9$ 0.17 0.001 0.61 0.24 $3.1\pm 1.0$ 0.12 $<10^{-6}$ 0.63 0.24
Tweeting
Highly influential ArXiv-first effect $1.4\pm 0.8$ 0.15 $2\times 10^{-4}$ 0.43 0.10 $0.06\pm 0.03$ 0.06 0.001 0.44 0.16
citations Revision Effect $2.1\pm 0.7$ 0.22 $<10^{-6}$ 0.44 0.10 $0.10\pm 0.01$ 0.10 $<10^{-6}$ 0.45 0.16
Effect of Tweeting $2.0\pm 0.4$ 0.18 $<10^{-6}$ 0.56 0.14 $0.05\pm 0.02$ 0.04 $<10^{-6}$ 0.47 0.19
Effect of Author(s) $2.6\pm 2.3$ 0.20 0.02 0.58 0.12 $0.20\pm 0.09$ 0.16 $<10^{-6}$ 0.54 0.17
Tweeting
Table 3: Average treatment effect (ATE) using standardization of each of our treatments on all citations and highly influential citations over 5 years after first publication. “Relative ATE” is calculated by dividing the ATE by the mean outcome value.
Treatment Year Computer Science Physics
TMLE ATE p-value TMLE ATE p-value
First $\mathbf{0.5\pm 0.5}$ $\mathbf{0.039}$ $\mathbf{0.37\pm 0.23}$ $\mathbf{0.001}$
Arxiv-first Second $\mathbf{2.7\pm 1.7}$ $\mathbf{0.001}$ $\mathbf{0.70\pm 0.23}$ $\mathbf{<10^{-6}}$
effect Third $\mathbf{4.7\pm 3.6}$ $\mathbf{0.010}$ $\mathbf{0.93\pm 0.27}$ $\mathbf{<10^{-6}}$
Fourth $6\pm 6.1$ $0.051$ $\mathbf{0.90\pm 0.24}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{7.2\pm 6}$ $\mathbf{0.017}$ $\mathbf{0.74\pm 0.29}$ $\mathbf{<10^{-6}}$
First $\mathbf{0.5\pm 0.4}$ $\bf{0.040}$ $\mathbf{0.71\pm 0.12}$ $\mathbf{<10^{-6}}$
Revision Second $\mathbf{2.5\pm 1.4}$ $\mathbf{0.008}$ $\mathbf{0.93\pm 0.17}$ $\mathbf{<10^{-6}}$
effect Third 4.5 $\pm$ 3.3 $\mathbf{0.008}$ $\mathbf{1.0\pm 0.19}$ $\mathbf{<10^{-6}}$
Fourth $5.6\pm 5.7$ $0.056$ $\mathbf{0.89\pm 0.20}$ $\mathbf{<10^{-6}}$
Fifth $4.9\pm 7.4$ $0.194$ $\mathbf{0.79\pm 0.20}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.5\pm 0.2}$ $\mathbf{<10^{-6}}$ $\mathbf{1.65\pm 1.03}$ $\mathbf{0.002}$
Effect of Second $\mathbf{4.9\pm 0.7}$ $\mathbf{<10^{-6}}$ $\mathbf{1.08\pm 0.63}$ $\mathbf{7\times 10^{-4}}$
tweeting Third $\mathbf{9.9\pm 1.6}$ $\mathbf{<10^{-6}}$ $\mathbf{1.90\pm 0.79}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{13.9\pm 2.5}$ $\mathbf{<10^{-6}}$ $\mathbf{1.52\pm 0.64}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{14.0\pm 3.2}$ $\mathbf{<10^{-6}}$ $\mathbf{1.54\pm 0.57}$ $\mathbf{<10^{-6}}$
First $\mathbf{1.1\pm 0.3}$ $\mathbf{<10^{-6}}$ $\mathbf{0.37\pm 0.10}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{1.1\pm 1}$ $\mathbf{0.037}$ ${-0.06\pm 0.62}$ 0.8540.854{0.854}0.854
author(s) Third $1.9\pm 2.03$ $0.072$ $\mathbf{0.55\pm 0.13}$ $\mathbf{<10^{-6}}$
tweeting Fourth $\mathbf{4.3\pm 3.2}$ $\mathbf{0.008}$ ${0.36\pm 1.12}$ $0.53$
Fifth $\mathbf{13.5\pm 5.5}$ $\mathbf{<10^{-6}}$ ${0.14\pm 0.39}$ $0.48$
@@
Treatment Year Computer Science Physics
TMLE ATE p-value TMLE ATE p-value
First $0.09\pm 0.10$ $0.108$ ${0.03\pm 0.03}$ 0.060.06{0.06}0.06
Arxiv-first Second $\mathbf{0.49\pm 0.36}$ $\mathbf{0.008}$ $\mathbf{0.07\pm 0.03}$ $\mathbf{<10^{-6}}$
effect Third $0.59\pm 0.90$ $0.196$ $\mathbf{0.08\pm 0.02}$ $\mathbf{<10^{-6}}$
Fourth $\mathbf{0.98\pm 0.96}$ $\mathbf{0.046}$ $\mathbf{0.06\pm 0.03}$ $\mathbf{2.5\times 10^{-5}}$
Fifth $\mathbf{1.10\pm 0.88}$ $\mathbf{0.015}$ $\mathbf{0.05\pm 0.03}$ $\mathbf{0.007}$
First $0.04\pm 0.10$ $0.426$ $\mathbf{0.03\pm 0.012}$ $\mathbf{<10^{-6}}$
Revision Second $\mathbf{0.46\pm 0.33}$ $\mathbf{0.005}$ $\mathbf{0.05\pm 0.02}$ $\mathbf{<10^{-6}}$
effect Third 0.73 $\pm$ 78 $0.063$ $\mathbf{0.05\pm 0.02}$ $\mathbf{<10^{-6}}$
Fourth $0.76\pm 1.28$ $0.244$ $\mathbf{0.04\pm 0.02}$ $\mathbf{<10^{-6}}$
Fifth $0.55\pm 1.60$ $0.501$ $\mathbf{0.03\pm 0.02}$ $\mathbf{5\times 10^{-5}}$
First $\mathbf{0.32\pm 0.05}$ $\mathbf{<10^{-6}}$ $\mathbf{0.09\pm 0.09}$ $\mathbf{0.05}$
Effect of Second $\mathbf{1.01\pm 0.16}$ $\mathbf{<10^{-6}}$ $\mathbf{0.08\pm 0.05}$ $\mathbf{0.002}$
tweeting Third $\mathbf{1.79\pm 0.33}$ $\mathbf{<10^{-6}}$ $\mathbf{0.09\pm 0.05}$ $\mathbf{7\times 10^{-5}}$
Fourth $\mathbf{2.20\pm 0.50}$ $\mathbf{<10^{-6}}$ $\mathbf{0.06\pm 0.06}$ $\mathbf{0.05}$
Fifth $\mathbf{2.11\pm 0.61}$ $\mathbf{<10^{-6}}$ $\mathbf{0.09\pm 0.06}$ $\mathbf{0.001}$
First $\mathbf{0.18\pm 0.05}$ $\mathbf{3.2\times 10^{-6}}$ $\mathbf{0.07\pm 0.02}$ $\mathbf{<10^{-6}}$
Effect of Second $\mathbf{0.63\pm 0.27}$ $\mathbf{6.8\times 10^{-6}}$ $\mathbf{0.03\pm 0.01}$ $\mathbf{<10^{-6}}$
author(s) Third $0.31\pm 0.47$ $0.187$ ${0.01\pm 0.03}$ 0.060.06{0.06}0.06
tweeting Fourth $\mathbf{2.43\pm 1.15}$ $\mathbf{3.2\times 10^{-5}}$ $\mathbf{0.10\pm 0.04}$ $\mathbf{<10^{-6}}$
Fifth $\mathbf{1.74\pm 0.80}$ $\mathbf{2\times 10^{-5}}$ $\mathbf{0.01\pm 0.01}$ $\mathbf{0.03}$
E. Average treatment effects using standardization
We find statistically significant positive effects for all our treatments for all citations and highly influential citations across computer science and physics. The results are listed in Table 3.
F. Average treatment effects for the first through fifth year of publication